OpenStudy (anonymous):

I'm struggling with an inverse laplace transform: 1/((s^2+1)^2) What I'd normally do is to use the convolution theorem, and get y(t)=sin^2(t), but I'm being a bit of a retard, and can't seem to get my answer to fit with the suggested solution.

OpenStudy (anonymous):

inverse laplace of \[1\div(s ^{2}+1)\] is sint using convolution we need to find \[\int\limits_{0}^{t} siny \sin(t-y) dy\]

OpenStudy (anonymous):

sinAsinB = 1/2[cos(A - B) - cos(A + B)]

OpenStudy (anonymous):

As I suspected, I'm a retard. I forgot to calculate the convolution after I had found the inverse transform. Thanks!

OpenStudy (anonymous):

welcome